OpenStudy (anonymous):

Please help!!! =( Elimination method 3x + 2y= 9 6x + 5y= 15

OpenStudy (angela210793):

step one: multiply the first equation by (-2) so u get -2*(3x+2y=9) ------> -6x-4y=-18 6x+5y=15 6x+5y=15 step two: add both equations ---------- 0+y=-3 y=-3 now that u have the value of y u can substitute that into one of the equations and find the value of x

OpenStudy (anonymous):

why is it -2?